# Novel High-Radiopurity Doped Amorphous Silicon Resistors for Low-Background Detectors

### v1

@@ 1 Introduction
We decided to investigate thin-film options and contacted Anritsu [2] in Morgan Hill CA. together with ViaMEMS Technologies [15], a fabrication business in Fremont CA. It was suggested that we investigate low pressure chemical vapor deposition (LPCVD) [8] polysilicon, but these firms were not positioned to initiate the required R&D. It is known that chemical vapor deposition of silicon is inherently an ultra-high purity process which is supported by the cleanroom procedures typically maintained at microfabrication facilities. Our own previous experience in the radioassay of materials for our experimental research had revealed that fused silica is available with sufficiently high radiopurity. Sapphire as a substrate for resistor fabrication was also expected to be suitable, although this was unconfirmed at the time. We decided to initiate our investigations “in house” at the Stanford Nanofabrication Facility (SNF). We obtained high purity Suprasil 300 fused silica tubing from Heraeus Inc. of the approximate dimensions needed (6 mm OD, 4 mm ID).
In our initial work, we collaborated with Maurice Stevens of Stanford Nanofabrication Facility (SNF) [11] who produced our first set of prototypes with undoped “intrinsic” silicon of 1.4 m thickness with satisfactory mechanical results but the resistance of these parts proved to be far too high, even at this large layer thickness. Our work continued at the Marvell Berkeley Nanofabrication Laboratory (MBNL) [9] with Sam Tsitrin to examine whether doping could be used to produce Si at our desired resistivity.
@@
The resistors double as structural members in the tensioned field cage.
@@
The resistors are cylindrical tubes with a resistive coated outer surface and with an ID to accommodate end-to-end tensioning rods for the field cage.
@@
These dual-purpose resistors, hereafter referred to as spacer/resistors, must have extremely low radioactive contamination, be cryogenically compatible at our operating temperature, and be mechanically robust under compressive loading. The optional requirement for reflectivity at 175 nm arises due to the need to collect as much of the expected VUV scintillation light produced in the LXe.
Our design studies concluded that the desired resistance per resistor in our voltage divider at our operating temperature of 165 K should be in the range of 0.1 G to 10.0 G, due to constraints on the maximum current draw from the HV power supply and the maximum acceptable RC discharge time constant. Acceptable field uniformity is achieved if the variation in the set of resistors within a divider chain is less than 20%.
Figure 1 shows the conceptual layout of the TPC field cage, with the field-shaping rings separated by 24 dielectric segments through which anode to cathode tensioning rods pass through holes in the rings and through the inner diameter of the spacers, to compress the cage. Some of the spacers (nominally 6 out of 24 rod segments) are dual-purpose spacer/resistors that make up the voltage divider. For the nominal nEXO field cage, these parts have length, outer diameter (OD) and inner diameter (ID) of approximately 17 mm, 6 mm and 4 mm respectively. The spacer/resistors have a resistive coating on their outer surface except at the ends which are metalized to make conductive contact with each field-shaping ring, thereby forming the voltage divider circuit. The absence of connectors/fasteners within the resistor chain and the insetting of the resistor ends from the edges of the field-shaping rings avoids the associated sharp points and high electric field regions, thereby reducing HV performance risk. The material used for these resistors (silicon thin film, fused silica or sapphire substrate, and the thin metallization) are not expected to produce chemical emanation.
@@ 3 Amorphous and Polycrystalline Silicon: Initial R&D at SNF
We were advised [1] that apparently a transition temperature existed between 580°C and 590°C, between the amorphous and polycrystalline states, with the amorphous state produced below 590°C. While we did not confirm the crystallinity of these different coatings by independent means, our observations convinced us that (plausibly) amorphous silicon (aSi) was more appropriate for our application.
@@
| Run 3 | , RMS >20% | matte, nonuniform | ||
| Run 4 | , RMS >50% | matte, nonuniform | ||
@@ 4 Experimental setup at SLAC
We were advised at SNF that the doping equipment available at the time was unlikely to reliably produce low dopant concentrations and were pointed to colleagues at Berkeley as an alternative. We then approached the Director of the Marvell Berkeley Nanofabrication Laboratory (MBNL), Dr. William Flounder, for assistance with our R&D program and a meeting was arranged.
@@ 5 Phosphorous Doped Silicon R&D at MBNL
At MBNL we devised a program using phosphorous doping with their Tystar model Mini-Tytan 3800 Horizontal LPCVD furnace [12]. A CHA Industries model Solution E-beam Evaporator [3] was used for metallization. Their system included a recently installed lower range gas flow controller (GFC) from Pivotal Systems [10], capable of flow rates as low as 0.2 standard cubic centimeters per minute (sccm). We used 150 mm silicon and fused silica wafers for trial resistivity runs, where a 200 sccm silane () flow rate was supplemented by varying flow rates of a dopant gas : a 50/50 mixture of silane and phosphine (). The furnace temperature was kept at 550°C for all runs to give us the best chance of depositing amorphous film at a reasonable deposition rate. Following the LPCVD process, metal contact strips (10 nm Ti undercoating with a 100 nm top Au layer) were evaporated perpendicular to the flat at both ends of each wafer. The wafers were then diced into 10 cm long x 0.5 cm wide rectangles (see Figure 4) which would fit in the testing fixture used for spacer/resistor prototypes. These diced planar resistors served as prototypes in resistivity testing to determine the needed doping recipe.
@@ 5.1 Low Temperature testing
After measuring the resistance of the wafer strips at room temperature, the testing was redone in the EC at 165 K as described in section 4. The total resistivity change from 295 K to 165 K due to TCR effects, a multiplicative factor we designate as , determines whether or not a particular doping level would produce a spacer/resistor within our target range of 0.1 - 1.0 G at the design operating temperature. Is was observed that a process using 0.7 sccm flow of the silane/phosphine mixture with a LPCVD layer thickness of about 300 nm achieved the required resistance range. The sheet resistance (measured using the 4-point contact method at MBNL) and aSi thickness (measured using ellipsometry, also at MBNL) are used to calculate the aSi resistivity. This resistivity is then used to predict the resistance of our standard spacer/resistor of dimensions (17 mm length, 6 mm OD, 4 mm ID) at room temperature, and using , also at 165 K.
@@ 5.2 Fabrication of spacer/resistors
This procedure currently accommodates 20 - 25 parts per run, see Figure 7. A purpose built fixture should be good for perhaps twice that number. The results for this production run are given in Table 5.2, showing the measured resistance at both room temperature and at the 165 K operating temperature. Figure 8 shows one spacer/resistor from the production run. The mirror-like aSi surface, the matte Pd-coated end, and the slight chamfer at the outer radius of the metallization are visible.
@@ 5.5 Comment on High Voltage testing
To date, no spacer/resistor has been tested at a potential difference above 1 kV end-to-end. Usually our resistance measurements with the Keithley electrometer are done at its 40 V auto setting, but tests at higher voltages up to 1 kV were made, and as expected for this modest voltage no breakdown was observed. For the nEXO design, with a nominal operating voltage at the cathode of 50 kV, the expected potential drop across each spacer/resistor is expected to be about 860 V. However, further HV testing to roughly 2 kV would be prudent. In the event issues were to be observed, the standard surface features for HV insulators designed to interrupt current paths such as circumferential ribs could be added. It has not yet been necessary to investigate this mitigation scheme by simulation or testing.
@@ 6 Radiopurity testing at PNNL
Ultra-high radiopurity is an important requirement for most components of the nEXO detector, determined by simulation to warrant a target radiopurity at the approximately sub-10 ppt level for U and Th for the spacer/resistors. The first prototypes made at SNF were sent to our radio-assay expert collaborators within the nEXO project, based at Pacific Northwest National Laboratory (PNNL). The PNNL team employs high-sensitivity inductively coupled plasma mass spectrometry (ICPMS) techniques to measure the U and Th content of some materials at the sub-ppt level [22, 18, 17]. We delivered a prototype completed spacer/resistor made at SNF along with an untreated silica tube blank to PNNL for analysis.
@@ 7 UV Reflectivity Studies at UMASS
A 4” witness wafer from our spacer/resistor run at SNF (aSi deposited at 550°C) was sent to our colleagues at UMASS who have a dedicated LXe reflectivity testing setup for a reflectivity measurement [13]. The results they obtained are given in Table 7, which indicate useful reflectivity at 175 nm and in principle improves light collection. Additional data and analysis will be required before both the reflectance and the specular and diffuse components can be determined [14]. As one can see from the TPC rendering in Figure 1, for the case of the nEXO design the spacers represent a small solid angle for light reflection compared to other surfaces. However, for some applications of this aSi technique to a larger surface area, the importance of reflectivity might be more significant.
@@ Acknowledgments
We are grateful for the contributions of Maurice Stevens, formerly of Stanford Nanofabrication Laboratory, whose initial participation in our R&D program was essential to get this project started, and to Dr.Mary Tang (Managing Director of SNF) for her support. We thank Dr.William Flounder, Director of the Marvell Berkeley Nanofabrication Laboratory (MBNL). We also appreciate the insights and advice we received from Prof.Ted Kamins (Stanford). Finally, we thank our colleagues within the nEXO Collaboration for input and discussions during this research.

### v2

@@ 1 Introduction
We decided to investigate thin-film options and contacted two industrial firms [3, 18] that suggested that we investigate low pressure chemical vapor deposition (LPCVD) [26] polysilicon, but these firms were not positioned to initiate the required R&D. It is known that chemical vapor deposition of silicon is inherently an ultra-high purity process which is supported by the cleanroom procedures typically maintained at microfabrication facilities. Our own previous experience in the radioassay of materials for our experimental research had revealed that fused silica is available with sufficiently high radiopurity. Sapphire as a substrate for resistor fabrication was also expected to be suitable, although this was unconfirmed at the time. We decided to initiate our investigations in house at the Stanford Nanofabrication Facility (SNF) [14]. We obtained high purity Suprasil 300 fused silica tubing from Heraeus Inc. of the approximate dimensions needed (6 mm OD, 4 mm ID).
At SNF we produced our first set of prototypes with undoped “intrinsic” silicon of 1.4 m thickness with satisfactory mechanical results but the resistance of these parts proved to be far too high, even at this large layer thickness. Our work continued at the Marvell Berkeley Nanofabrication Laboratory (MBNL) [11] to examine whether doping could be used to produce Si at our desired resistivity.
@@
The resistors double as structural members in the tensioned field cage and must be mechanically robust under compressive loading .
@@
The resistors are cylindrical tubes with a resistive coated outer surface and with an inner diameter to accommodate end-to-end tensioning rods for the field cage.
@@
These dual-purpose resistors, hereafter referred to as spacer/resistors, are required to have the properties listed above except for the last one listed. The desirability of reflectivity at 175 nm arises due to the need to collect as much of the expected VUV scintillation light produced in the LXe, a need that depends on the total solid angle reflected by these parts.
Our design studies concluded that the desired resistance per resistor in our voltage divider at our operating temperature of 165 K should be in the range of 0.1 G to 10.0 G, due to constraints on the maximum current draw from the HV power supply and the maximum acceptable RC discharge time constant (The estimated capacitance of the completed TPC field cage is 3 nF, corresponding to a few minutes for a maximum total resistance of the voltage divider of approximately 100 GOhm). Acceptable field uniformity is achieved if the variation in the set of resistors within a divider chain is less than 20%.
Figure 1 shows the conceptual layout of the TPC field cage, with the field-shaping rings separated by 24 dielectric spacers through which anode to cathode tensioning rods pass through holes in the rings and through the inner diameter of the spacers, to compress the cage. Some of the spacers (nominally 6 out of 24) are dual-purpose spacer/resistors that make up the voltage divider. For the nominal nEXO field cage, these parts have length, outer diameter (OD) and inner diameter (ID) of approximately 17 mm, 6 mm and 4 mm respectively. The spacer/resistors have a resistive coating on their outer surface except at the ends which are metalized to make conductive contact with each field-shaping ring, thereby forming the voltage divider circuit. The absence of connectors/fasteners within the resistor chain and the insetting of the resistor ends from the edges of the field-shaping rings avoids the associated sharp points and high electric field regions, thereby reducing HV performance risk. The material used for these resistors (silicon thin film, fused silica or sapphire substrate, and the thin metallization) are not expected to produce chemical emanation.
@@ 3 Amorphous and Polycrystalline Silicon: Initial R&D at SNF
We were advised [2] that apparently a transition temperature existed between 580°C and 590°C between the amorphous and polycrystalline states of Si, with the amorphous state produced below 590°C, and that a precise transition temperature value was not found in the literature. While we did not confirm the crystallinity of these different coatings by independent means, our observations convinced us that (plausibly) amorphous silicon (aSi) was more appropriate for our application.
@@
| Run 3 | , SD >20% | matte, nonuniform | ||
| Run 4 | , SD >50% | matte, nonuniform | ||
@@ 4 Experimental setup at SLAC
We were advised at SNF that the doping equipment available at the time was unlikely to reliably produce low dopant concentrations and were pointed to colleagues at Berkeley as an alternative. We then approached the Director of the Marvell Berkeley Nanofabrication Laboratory (MBNL) for assistance with our R&D program.
@@ 5 Phosphorous Doped Silicon R&D at MBNL
At MBNL we devised a program using phosphorous doping with their Tystar model Mini-Tytan 3800 Horizontal LPCVD furnace [15]. Their system included a recently installed lower range gas flow controller (GFC) from Pivotal Systems [13], capable of flow rates as low as 0.2 standard cubic centimeters per minute (sccm). A CHA Industries model Solution E-beam Evaporator [5] was used for metallization. We used 150 mm silicon and fused silica wafers for trial resistivity runs, where a 200 sccm silane () flow rate was supplemented by varying flow rates of a dopant gas : a 50/50 mixture of silane and phosphine (). The furnace temperature was kept at 550°C for all runs to give us the best chance of depositing amorphous film at a reasonable deposition rate. Following the LPCVD process, metal contact strips (10 nm Ti undercoating with a 100 nm top Au layer) were evaporated perpendicular to the flat at both ends of each wafer. The wafers were then diced into 10 cm long x 0.5 cm wide rectangles (see Figure 4) which would fit in the testing fixture used for spacer/resistor prototypes. These diced planar resistors served as prototypes in resistivity testing to determine the needed doping recipe.
@@
The procedures we use to characterize our produced aSi parts are common to all the results we obtained at MBNL and are summarized as follows :
We used standard nano-fabrication methods to determine the properties of our aSi films. Sheet resistance is provided by four-point resistance mapping instruments [4], spectroscopic ellipsometry and reflectometry [7, 12] as well as a mechanical technique known as profilometry [1] give us the film thickness. The precision of the data for sheet resistance and thickness of each film sample arises from instrumental effects that are estimated and reported by each device on a measurement-by-measurement basis. The uncertainties in the physical results are propagated into the calculation of derived quantities of interest such as the film resistivity or the expected resistance of a fabricated spacer/resistor. If necessary, “witness wafers”, both silicon (with an oxide layer) or silica, are used as control substrates and are included in the furnace during the aSi CVD coating process. For sheet resistance and ellipsometry measurements, multiple points on each sample or witness wafer surface are measured and the observed variation contributes to the reported measurement uncertainty.
@@ 5.1 Low Temperature testing
After measuring the resistance of the wafer strips at room temperature, the testing was redone in the EC at 165 K as described in section 4. The total resistivity change from 295 K to 165 K due to TCR effects, a multiplicative factor we designate as , determines whether or not a particular doping level would produce a spacer/resistor within our target range of 0.1 - 10.0 G at the design operating temperature. Is was observed that a process using 0.7 sccm flow of the silane/phosphine mixture with a LPCVD layer thickness of about 300 nm achieved the required resistance range. The sheet resistance (measured using the 4-point contact method at MBNL) and aSi thickness (measured using ellipsometry, also at MBNL) are used to calculate the aSi resistivity. This resistivity is then used to predict the resistance of our standard spacer/resistor of dimensions (17 mm length, 6 mm OD, 4 mm ID) at room temperature, and using , also at 165 K.
@@ 5.2 Fabrication of spacer/resistors
This procedure currently accommodates 20 - 25 parts per run, see Figure 7. A purpose built fixture should be good for perhaps twice that number. The results for this production run are given in Table 5.2, showing the measured resistance at both room temperature and at the 165 K operating temperature. Figure 8 shows one spacer/resistor from the production run. The mirror-like aSi surface and the matte Pd-coated end are visible. There is a slight chamfer at the outer radius of the metallization, which is only 114 microns wide (see Figure 5).
@@ 5.5 Comment on High Voltage testing
To date, no spacer/resistor has been tested at a potential difference above 1 kV end-to-end. Usually, our resistance measurements with the Keithley electrometer are done at its 40 V auto setting, but tests at higher voltages up to 1 kV were made, and as expected for this modest voltage no breakdown was observed. For the nEXO design, with a nominal operating voltage at the cathode of 50 kV, the expected potential drop across each spacer/resistor is expected to be about 860 V. However, further HV testing to roughly 2 kV would be prudent. In the event issues were to be observed, the standard surface features for HV insulators designed to interrupt current paths such as circumferential ribs could be added. It has not yet been necessary to investigate this mitigation scheme by simulation or testing.
Issues regarding the Ohmic behavior of these resistors have not yet been carefully studied as the original application envisioned in the nEXO TPC does not require good linearity. It has been observed that the large TCR does introduce small non-linearity due to dissipation heat at high current operation up to 1 kV. This effect should be mitigated by TPC cryostat temperature regulation.
@@ 6 Radiopurity testing at PNNL
Ultra-high radiopurity is an important requirement for most components of the nEXO detector, determined by simulation to warrant a target radiopurity at the approximately sub-10 ppt level for U and Th for the spacer/resistors. The first prototypes made at SNF were sent to our radio-assay expert collaborators within the nEXO project, based at Pacific Northwest National Laboratory (PNNL). The PNNL team employs high-sensitivity inductively coupled plasma mass spectrometry (ICPMS) techniques to measure the U and Th content of some materials at the sub-ppt level [25, 24, 21, 20]. We delivered a prototype completed spacer/resistor made at SNF along with an untreated silica tube blank to PNNL for analysis.
Concentrations are reported in ppt or pg/g for 232Th and 238U, and ppb or ng/g for natural K. One sample of each material was measured after full dissolution. The uncertainty on the measurement is the analytical precision from the instrumental analysis, which is the standard deviation from three acquisitions of the sample solution on the ICP-MS.
@@ 7 UV Reflectivity Studies at UMASS
A 4” witness wafer from our spacer/resistor run at SNF (aSi deposited at 550°C) was sent to our colleagues at UMASS who have a dedicated LXe reflectivity testing setup for a reflectivity measurement [16]. The results they obtained are given in Table 7 and agree within errors with the value available in the literature [27] which indicate useful reflectivity at 175 nm and in principle improves light collection. Additional data and analysis will be required before both the reflectance and the specular and diffuse components can be determined [17]. As one can see from the TPC rendering in Figure 1, for the case of the nEXO design the spacers represent a small solid angle for light reflection compared to other surfaces. However, for some applications of this aSi technique to a larger surface area, the importance of reflectivity might be more significant.
@@ Acknowledgments
We are grateful for the contributions of Maurice Stevens, formerly of Stanford Nanofabrication Laboratory, whose initial participation in our R&D program was essential to get this project started, and to Dr.Mary Tang (Managing Director of SNF) for her support. We thank Dr.William Flounders, Director of the Marvell Berkeley Nanofabrication Laboratory (MBNL). We also appreciate the insights and advice we received from Prof.Ted Kamins (Stanford). We thank Drs.Christopher Kenney and Julie Segal (SLAC Instrumentation Division) for their expertise and support. Finally, we thank our colleagues within the nEXO Collaboration for input and discussions during this research.